OpenStudy (anonymous):

What is the integral of 1/x(lnx)^3 dx

OpenStudy (anonymous):

Set up the integral and since you have an LN in the integral you see that you may have to use substitution or integration by parts

OpenStudy (anonymous):

∫ 1/{x * [ln(x)]^3} dx = ∫ 1/x * 1/[ln(x)]^3 dx. So let u = ln(x) <==> du = 1/x dx. Then, applying these substitutions yields: ∫ 1/x * 1/[ln(x)]^3 dx = ∫ 1/u^3 du = ∫ u^(-3) du = u^(-3 + 1)/(-3 + 1) + C = -1/(2u^2) + C = -1/{2 * [ln(x)]^2} + C

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 1 }{ x } (\ln x )^{3} \]

OpenStudy (anonymous):

\[ \int \frac{(\ln x)^3}{x}dx = \int (\ln x)^3d(\ln x) \]

OpenStudy (anonymous):

actually the (lnx)^3 is on the bottom xD

OpenStudy (anonymous):

thats what i did..

OpenStudy (anonymous):

(inx)^3 as you see is listed below.

OpenStudy (anonymous):

thank-you, I was telling wio :)

OpenStudy (anonymous):

Oh okay! your welcome. Goodluck!!!

OpenStudy (anonymous):

Thanks!!